CLICAL version 4, August 1992 [release September 1994]

User Manual Appendix

Changes made to the manual of version 2

Reversion x̃ of x is typed as x~ Conjugation x̄ of x is typed as x~' [in the manual typed as x~ and written as x̄]

New functions have been added: > Q(x) = ⟨xx̃⟩_{0} quadratic form of x [in the manual quad(x)] > N(x) = ⟨xx̄⟩_{0} negative quadratic form of x

A command has been altered: > FIX k displays outputs with k decimal places of precision [k = 3, …, 9]

New external files have been added. Try the following:

> GET GUIDE
> GET DEMO
> GET SPINOR
> GET TRIALITY

The command INFO has been extended so that you can see also the automorhism groups of scalar product of spinors.